OpenStudy (anonymous):

Suppose the integral from 2 to 8 of g of x, dx equals 13, and the integral from 6 to 8 of g of x, dx equals negative 3, find the value of 2 plus the integral from 2 to 6 of g of x, dx.

OpenStudy (anonymous):

https://gyazo.com/41bc06711b55841c5edc26d60f3d956a

OpenStudy (freckles):

\[\int\limits_2^6g(x) dx=\int\limits_2^8 g(x) dx- \int\limits_6^8 g(x)dx\]

OpenStudy (freckles):

we have from 2 to 8 given but we wanted 2 to 6 so we have to subtract the part out that is from 6 to 8

OpenStudy (freckles):

which is given also

OpenStudy (freckles):

don't forget to add the 2 in

OpenStudy (anonymous):

So how would go about to solve? Solve each integral then subtract?

OpenStudy (anonymous):

no just subtract

OpenStudy (anonymous):

Oooooooooohhhhhhh I'm an idiot, read the question so wrong. 16+2=18 Final answer 18 right?

OpenStudy (freckles):

right you have 2+13-(-3) which is 15+3 which is 18